### 06 May 2026 v1

# UniPCB: A Generation-Assisted Detection Framework for PCB Defect Inspection

Abstract
Printed Circuit Board (PCB) defect inspection faces two compounding challenges: scarce and imbalanced defect samples that limit model training, and insufficient feature representation under complex circuit backgrounds. Existing generation methods rely on single-modality conditions with coarse structural control, while detection methods improve architectures without addressing the data bottleneck. To resolve both challenges jointly, we propose a generation-assisted PCB defect inspection framework that integrates controlled defect synthesis with task-specific defect detection. On the generation side, a Multi-modal Condition Generator extracts complementary edge, depth, and text conditions in parallel. A ScaleEncoder then embeds these conditions into the diffusion U-Net at four resolutions, and a Condition Modulation applies FiLM-style spatially-adaptive modulation at each scale, enabling structurally aligned and defect-aware sample synthesis. On the detection side, an Inverted Residual Shift Attention couples self-attention with shift-wise convolution to jointly capture global context and local texture, and a Cross-level Complementary Fusion Block generates pixel-level gates for selective cross-level feature fusion. The synthesized samples directly enrich the detection training set, so that improvements in generation compound with improvements in detection. Extensive experiments on DsPCBSD+ demonstrate that UniPCB achieves mAP@0.5 of 98.0% and mAP@0.5:0.95 of 61.8% on defect detection, surpassing all compared methods, while the generation branch attains an FID of 129.61 and SSIM of 0.619, outperforming existing conditional generation approaches.
Index Terms:
PCB defect detection, Diffusion model, Multi-modal data generation.
I Introduction
Printed circuit boards (PCBs) are essential to modern electronics, but complex manufacturing processes can introduce defects such as shorts, opens, and spurious copper that threaten product reliability. Automated optical inspection (AOI) systems typically rely on reference-image comparison and handcrafted feature matching, making them sensitive to lighting variation, board layout changes, and novel defect morphologies, which can lead to high false-negative rates in production environments [69, 14].
Deep learning-based detectors offer substantially greater robustness by learning discriminative representations directly from data. However, their effectiveness depends on the availability of large volumes of labeled defect samples, which is fundamentally difficult to satisfy in PCB manufacturing. Defective boards are rare by design, and defect categories often follow a severely long-tailed distribution. As a result, models trained on limited available data tend to underfit rare defect categories and generalize poorly to unseen board layouts.
To alleviate this data bottleneck, traditional augmentation techniques such as rotation and flipping have been widely adopted, but they fail to capture the complex and variable defect patterns encountered in real production environments. Deep generative models offer a more principled alternative: VAEs [25, 51] and GANs [11] have been applied to defect synthesis, but VAEs tend to produce blurry images due to pixel-wise reconstruction losses, and GANs suffer from training instability and mode collapse. Diffusion models [16, 43, 44, 66, 42, 65] overcome these limitations with superior sample quality and diversity. Nonetheless, their training on natural image datasets limits their effectiveness on industrial images such as PCBs [5, 13], and existing PCB-oriented methods typically rely on a single control condition injected at the initial resolution, providing insufficient structural guidance for the dense, fine-grained circuit patterns unique to PCB images.
Figure 1: Comparison between UniPCB and previous approaches for PCB defect detection. (a) Previous methods address the two challenges in isolation: generation methods rely on single-modality conditions with limited structural control, while detection methods improve architectures but remain bottlenecked by scarce training data. (b) UniPCB jointly addresses both within a unified framework: the generation branch synthesizes high-quality defect samples while the detection branch exploits the enriched data through task-specific feature modeling modules.
Beyond high-quality sample generation, the expressive capability of the detection framework itself is equally critical. DETR [4] leverages end-to-end global modeling to significantly enhance defect perception in complex scenes, but its slow convergence and high computational cost hinder real-time industrial deployment. RT-DETR [68] improves inference speed while maintaining competitive accuracy through optimized attention mechanisms and detection head design, making it a more practical alternative. However, under complex backgrounds or small defects, its feature modeling and multi-scale feature fusion remain insufficient. The challenge of simultaneously optimizing detection accuracy and computational efficiency has been extensively studied in the general detection literature [47], yet achieving this balance for fine-grained PCB defects under dense circuit patterns remains an open problem [50].
These two issues often appear together in PCB inspection and jointly limit practical detection performance. On the one hand, a detection model trained on scarce and imbalanced data cannot fully leverage improvements in model architecture, as the data bottleneck fundamentally limits what the model can learn. On the other hand, a generation model producing low-fidelity or poorly controlled defect samples provides little benefit to detection, even when the detector itself is powerful. This coupling implies that addressing either challenge in isolation yields diminishing returns: the two must be resolved jointly within a coherent framework to achieve meaningful gains in PCB defect inspection.
To break this mutual bottleneck, we propose a unified generative–detection framework, termed UniPCB, which simultaneously enhances data quality through controlled defect synthesis and improves detection capability through task-specific architectural design. Fig. 1 compares UniPCB with previous approaches. As illustrated in Fig. 1(a), prior work treats the two problems separately: diffusion-based methods reconstruct images to identify pixel-level discrepancies but cannot generate new training data, while object detection methods improve architectures but remain fundamentally bottlenecked by scarce labeled data. As shown in Fig. 1(b), UniPCB unifies both within a single framework. On the generation side, a multi-modal conditional strategy synthesizes diverse and realistic defect samples by injecting edge, depth, and text conditions at multiple scales, ensuring structural alignment and semantic fidelity to real PCB images. On the detection side, the generated data directly serves as augmented training input. To further strengthen the detector’s sensitivity to small, low-contrast defects under complex circuit backgrounds, two task-specific modules are introduced: the IRSA Block for joint global–local feature modeling, and the CLCF Block for cross-level selective feature fusion. Through this co-design of generation and detection, each component amplifies the benefit of the other.
In summary, our main contributions are as follows.
❑  We propose UniPCB, a generation-assisted PCB defect inspection pipeline that couples controlled defect synthesis with a task-specific detector. Rather than treating data augmentation and detector design as isolated components, UniPCB evaluates how structured synthetic samples and PCB-oriented feature modeling jointly improve defect inspection performance.
❑  On the generation side, we design a Multi-modal Condition Generator that constructs parallel edge, depth, and text conditions, providing diverse and structured control signals to alleviate data scarcity and class imbalance. Building on this, we propose a latent diffusion-based synthesis network with a ScaleEncoder that embeds conditions at four resolutions and a CondMod block that jointly modulates noise and textual features, improving sample fidelity and structural consistency.
❑  On the detection side, we propose an Inverted Residual Shift Attention Block that couples self-attention and shift-wise convolution in cascaded within an inverted-residual structure, effectively capturing global context and local texture cues for small, low-contrast PCB defects.
II Related Work
II-A PCB Defect Generation
In PCB defect detection, deep learning models require sufficient labeled data, yet real defect samples are scarce, imbalanced, and costly to annotate. Synthetic defect generation has therefore been explored as an effective way to enrich training data and improve detector generalization. Existing methods mainly follow two technical routes: GAN-based synthesis and diffusion-based conditional generation.
Early defect generation methods often adopted GAN-based frameworks to synthesize additional defective samples for data augmentation [9, 10, 29]. By learning the distribution of defect textures and backgrounds, these methods can increase sample diversity under limited-data settings. Related conditional generation techniques, such as self-attention based generation [61] and spatially-adaptive normalization [38], further show the importance of long-range dependency modeling and spatially-aware modulation for preserving structural consistency. However, GAN-based methods are generally limited by unstable training, mode collapse, and insufficient control over the spatial relationship between defects and complex PCB backgrounds.
Recently, diffusion models have become a more promising alternative because of their stable training process, strong texture modeling ability, and flexible conditional control [49, 19]. In PCB-related defect generation, Deng et al. [7] introduced feature modulation and normalized flow modules to coordinate local and non-local information for learning high-quality defect distributions. Beyond PCB-specific settings, controllable diffusion methods such as T2I-Adapter [35] and Uni-ControlNet [41] demonstrate that external structural conditions can effectively guide pretrained diffusion models and improve controllability. These studies suggest that diffusion-based generation is well suited for synthesizing fine-grained industrial defect images.
Despite these advances, existing defect generation methods still have two limitations. First, most methods rely on limited or single-form conditions, which are insufficient to describe the dense circuit structures and subtle defect patterns in PCB images. Second, conditional signals are often injected in a coarse or single-scale manner, making it difficult to preserve fine traces and local defect geometry across the denoising process. To address these issues, we introduce a multi-modal condition generator together with multi-scale condition embedding and modulation, enabling more structured and defect-aware PCB sample synthesis for downstream detection.
II-B PCB Defect Detection
PCB defect detection differs from natural image detection because PCB defects are usually small, sparse, irregularly shaped, and visually similar to normal circuit patterns. Traditional handcrafted or template-matching methods are sensitive to manufacturing variations and therefore generalize poorly across board layouts and imaging conditions. Deep learning based detectors have improved robustness by learning discriminative representations automatically, and recent studies [60, 54, 59, 26, 32, 53, 34] mainly improve PCB defect detection from two perspectives: adaptive feature modeling and multi-scale feature fusion.
Adaptive feature modeling [31, 28, 12] aims to enhance defect-relevant responses while suppressing complex circuit background interference. Liu et al. [31] introduced a multi-head nonlocal transformer module to capture long-range dependencies and focus on defect regions. Guo et al. [12] proposed an Image-to-Instance Contextual Enhancement Component that incorporates global contextual information to reweight proposal features. Dai et al. [6] designed Dynamic Head, which dynamically modulates detection features with scale-aware, spatial-aware, and task-aware attention. These methods demonstrate the importance of adaptive representation learning for subtle defect perception.
Another line of work focuses on multi-scale feature fusion [56, 3, 22], which is critical for localizing tiny defects while preserving semantic discrimination. Yang et al. [56] designed a Coordinate Feature Refinement module to fuse channel- and spatial-level features and improve scale robustness. Cao et al. [3] proposed MRC-DETR, which embeds multi-scale residual coupling into a DETR-based framework for bare-board PCB defect detection. Ji et al. [22] introduced a Slim-Scale Adaptive Fusion structure that aggregates hierarchical features using grouped spatial convolutional pyramids and channel shuffle, improving fusion efficiency and detection accuracy.
Although these methods improve feature representation and localization, they mainly operate at the detector architecture level and typically assume that the training distribution is sufficiently representative. In practical PCB inspection, however, defect samples remain scarce and highly imbalanced, especially for rare defect categories. As a result, architecture-level improvements alone may still be limited by insufficient training data. This motivates our generation-assisted detection pipeline, which combines structured defect synthesis with task-specific feature modeling to address both data scarcity and representation limitations.
Figure 2: Overview of the proposed defect generation framework. Real PCB images captured by an AOI machine are processed by the Multi-modal Condition Generator (MmCG) to extract edge, depth, and text conditions, which are embedded into the Stable Diffusion model via ScaleEncoder and CondMod to synthesize defect samples. The generated images are merged with real data to form the mixed training dataset.
III Method
III-A Overview architecture
III-A1 Generation Framework
As illustrated in Fig. 2, the generation network is built upon Stable Diffusion [42] with ControlNet [63] and extended with three dedicated modules for non-natural industrial images. A real PCB image first passes through the Multi-modal Condition Generator (MmCG), which produces complementary edge, depth, and text conditions in parallel. These conditions are then embedded into the diffusion U-Net at four resolutions by the ScaleEncoder, and at each resolution a CondMod block jointly modulates the noise feature with the condition and textual embedding to steer the denoising process. The synthesized defect images are merged with real data to form a balanced training set for downstream detection. The three modules are detailed in Sec. III-B.
III-A2 Detection Framework
As shown in Fig. 4, the detection network extends RT-DETR [68] with a ResNet-18 backbone and two task-specific modules. IRSA Blocks replace the original BasicBlocks at every backbone stage to produce more discriminative features, and three CLCF blocks (CLCFL, CLCFM, CLCFS) subsequently fuse the backbone features with the CCFM outputs at pyramid levels S3, S4, and S5. The fused features are concatenated and fed to the IoU-aware query selection and decoder. The two modules are detailed in Sec. III-C.
Figure 3: (a) Structure of the ScaleEncoder, which encodes the condition map via ScaleAlign layer and injects multi-resolution features into the diffusion U-Net via ScaleEmbed layer. (b) Structure of the Condition Modulation (CondMod), which applies FiLM-style spatially-adaptive modulation to fuse condition features and textual embeddings with the noise map at each resolution.
III-B Generation Framework
III-B1 Latent Diffusion Model
We build our generator on the Latent Diffusion Model (LDM) [42], which performs denoising in the latent space of a pretrained autoencoder rather than in pixel space, substantially reducing computational cost. Given an input image $x\in\mathbb{R}^{H\times W\times 3}$, the encoder $\mathcal{E}$ compresses it into a latent code $z=\mathcal{E}(x)$, on which forward diffusion and reverse denoising are performed following the standard DDPM formulation [16]:
$q(z_{t}\mid z_{0})=\mathcal{N}\!\left(z_{t};\,\sqrt{\bar{\alpha}_{t}}\,z_{0},\,(1-\bar{\alpha}_{t})\mathbf{I}\right),\quad\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i}),$ (1)
[…]
III-B2 Multi-modal Condition Generator
Diffusion models require paired (image, condition) data, yet no public image–text PCB dataset exists. To automate dataset construction and provide rich control signals, we design a Multi-modal Condition Generator (MmCG) with three parallel branches, illustrated in Fig. 2.
Edge branch. PCB traces and substrates differ sharply in color and brightness, making edge maps a natural structural prior. We employ the Canny operator [2] with high and low thresholds adaptively derived from the Otsu threshold [36] (scaled by fixed upper and lower factors), yielding image-specific edges without manual tuning.
Depth branch. Edge maps alone cannot distinguish foreground traces from background substrate, which may cause spatial misalignment. We therefore estimate a depth map using the pretrained Depth-Anything V2 [57], supplying complementary scene-layout cues.
Text branch. Since no public image–text PCB dataset exists, we construct structured prompts automatically from detection outputs. A pretrained YOLO11 [24] first predicts defect categories and bounding boxes on newly collected PCB images. Each defect is then characterized along three axes: scale, classified as small, medium, or large via a dual-threshold rule applied to the bounding-box area; location, encoded by mapping the normalized box-center coordinates onto a $3{\times}3$ spatial grid (top, bottom, left, right, center, and four corners); and distribution, described at the instance level when defect count is low (per-defect category, scale, and position) or abstracted to region-level patterns (e.g., scattered or locally clustered) when count is high. The resulting descriptions are slotted into a template library parameterized by {category, scale, location, quantity} to produce structured prompts, which are subsequently encoded by the frozen CLIP text encoder.
Together, the three branches provide structural, spatial, and semantic conditions that jointly guide the downstream diffusion process.
III-B3 ScaleEncoder
ControlNet [63] injects condition features only at the initial resolution, which is inadequate for PCB images with dense, fine-grained circuit patterns. Following the multi-scale injection principle of Uni-ControlNet [67], we propose a ScaleEncoder that embeds conditions into the diffusion U-Net at four resolutions ($\times64$, $\times32$, $\times16$, $\times8$).
As shown in Fig. 3(a), ScaleEncoder consists of two components. ScaleAlign encodes the raw condition map through three SiLU-activated standard convolutions interleaved with two SPD-Conv downsampling layers [46]. SPD-Conv reorganizes spatial information into the channel dimension, preserving local structural cues that standard strided convolutions tend to lose, a property well suited to thin PCB traces. ScaleEmbed then taps intermediate features at four resolutions and refines each via a DWConv–SiLU block followed by a zero convolution [63], whose zero initialization allows the pretrained diffusion backbone to remain undisturbed at the start of training and gradually learn the optimal embedding. The four refined features are fed to the corresponding Condition Modulation.
III-B4 Condition Modulation
To strengthen condition guidance at each resolution, we introduce a Condition Modulation (CondMod), shown in Fig. 3(b). At every scale, parallel CondMod fuse the noise feature, the condition feature, and the textual embedding.
Following the FiLM-style modulation [40, 8], the noise feature is first normalized via GroupNorm to obtain $\mathcal{N}_{norm}$, and the condition feature is passed through two parallel depth-wise convolutions to predict the scale and shift parameters $\alpha$ and $\beta$:
$\mathbf{X}_{\mathrm{mid}}=\mathcal{N}_{norm}\odot(1+\alpha)+\beta,$ (2)
where $\odot$ denotes element-wise multiplication. The $(1+\alpha)$ formulation initializes the block close to an identity mapping, stabilizing training. The modulated feature $\mathbf{X}_{\mathrm{mid}}$ is then added to the textual embedding to yield the block output, achieving spatially-aware conditioning driven by the visual control signal and semantically-aware conditioning driven by the prompt.
Figure 4: Overview of the proposed detection framework. IRSA Blocks replace BasicBlocks in all four stages of the ResNet-18 backbone. Yellow and blue blocks in the Neck denote upsampling and downsampling operations, respectively.
III-C Detection Framework
III-C1 Inverted Residual Shift Attention
PCB defects are typically small, low-contrast, and easily masked by complex circuit patterns. Self-attention captures global semantics but is insensitive to fine textures, while convolutions capture local details but lack long-range context; neither alone is sufficient for reliable small-defect detection.
We therefore propose the Inverted Residual Shift Attention (IRSA) block (Fig. 5(a)), which couples self-attention and shift-wise convolution in cascaded within an inverted-residual structure. Given input $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, we first apply a convolution–batch normalization–ReLU (CBR) layer to obtain a preliminary feature $\mathbf{X}_{\text{pre}}=\mathrm{CBR}(\mathbf{X})$, and then expand the channel dimension from $C$ to $C^{\prime}$ via a $1{\times}1$ convolution, yielding $\mathbf{X}_{\text{expand}}$. Two cascaded modules are then employed:
$\mathbf{X}_{{att}}={Softmax}\!\left(\frac{\mathbf{Q}\mathbf{K}^{\!\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$ (3)
$\mathbf{X}_{{swc}}={DWConv}\!\left({Shift}_{G}(\mathbf{X}_{{att}})\right),$ (4)
where $\mathbf{Q}$ and $\mathbf{K}$ are linear projections of $\mathbf{X}_{{pre}}$, and $\mathbf{V}$ is a linear projections of $\mathbf{X}_{{expand}}$. ${Shift}_{G}(\cdot)$ splits channels into $G$ groups and spatially shifts each group along one of $G$ pre-defined directions (up, down, left, right, and four diagonals) [27], followed by a depth-wise convolution that aggregates the shifted context. The branch outputs are merged and projected back to $C$ channels:
$\mathbf{X}_{{out}}={Conv}_{1\times 1}(\mathbf{X}_{{att}}+\mathbf{X}_{{swc}}).$ (5)
The final output restores the input and the preliminary feature paths to ensure stable gradient flow:
$\mathbf{X}_{{final}}=\mathbf{X}+\mathbf{X}_{{pre}}+\mathbf{X}_{{out}}.$ (6)
The cascaded coupling of global attention and shift-wise convolution, together with the multi-level residuals, yields efficient yet discriminative features for small PCB defects.
Figure 5: (a) Overview of the Inverted Residual Shift Attention Block. (b) Structure of the shift-wise convolution (SWC).
III-C2 Cross-level Complementary Fusion
Shallow features retain fine textures crucial for localizing tiny defects but are noise-sensitive, whereas deep features carry strong semantics but lose spatial detail. Naive addition or concatenation tends to amplify noise or suppress subtle defects. To better fuse the two, we propose the Cross-level Complementary Fusion (CLCF), comprising a Dual-path Collaborative Attention (DPCA) that predicts a pixel-wise gate, and a gated weighting scheme with residual compensation (Fig. 6(a)).
DPCA sub-block. Given the concatenated input $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, a $1{\times}1$ convolution produces the shared tensor $\mathbf{X}_{qkv}$. Two parallel branches then operate on it. The local branch captures fine-grained cues:
$\mathbf{X}_{\text{loc}}=\mathrm{GConv}\!\left(\mathrm{CS}\!\left(\mathrm{DWConv}_{3\times 3}(\mathbf{X}_{qkv})\right)\right),$ (7)
where $\mathrm{CS}(\cdot)$ denotes channel shuffle. The global branch computes self-attention with a learnable per-head scaling vector $\bm{\alpha}$, following [1]:
$\mathbf{X}_{\text{glo}}=\mathrm{Conv}_{1\times 1}\!\left(\mathrm{Softmax}\!\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\bm{\alpha}}\right)\mathbf{V}\right)+\mathbf{X}_{qkv}.$ (8)
The two outputs are concatenated with the input and passed through a grouped convolution with sigmoid activation to yield a pixel-wise gate:
$\mathbf{w}=\sigma\!\left(\mathrm{GConv}_{7\times 7}\!\left(\mathrm{CS}\!\left(\mathrm{Concat}[\mathbf{X}_{\text{loc}},\mathbf{X}_{\text{glo}},\mathbf{X}]\right)\right)\right),$ (9)
Gated fusion. Let $\mathbf{F}_{l}$ and $\mathbf{F}_{h}$ denote the low- and high-level inputs aligned to the same resolution. CLCF fuses them via a gate-and-residual scheme:
$\mathbf{F}_{f}=\mathrm{Proj}\!\left(\mathbf{w}\odot\mathbf{F}_{l}+(1-\mathbf{w})\odot\mathbf{F}_{h}+(\mathbf{F}_{l}+\mathbf{F}_{h})\right),$ (10)
where the residual term $\mathbf{F}_{l}+\mathbf{F}_{h}$ prevents information loss from over-gating. The gate $\mathbf{w}$ adaptively selects detail cues from $\mathbf{F}_{l}$ where defects are subtle and semantic cues from $\mathbf{F}_{h}$ elsewhere, yielding features well suited for the detection head.
Figure 6: (a) Overview of the Cross-level Complementary Fusion (CLCF) Block. (b) Structure of the Dual-path Collaborative Attention (DPCA) Block.
Figure 7: Representative annotated samples from the six defect categories, including single-class images (a–f) and a multi-class example (g). (a) Short; (b) Spur; (c) Spurious copper; (d) Open; (e) Mouse bite; (f) Hole breakout; (g) Multiple classes.
IV Experiments
IV-A Datasets
All experiments are based on the public PCB defect dataset DsPCBSD+ [33]. Six defect categories that significantly affect PCB quality are selected: short, spur, spurious copper, open, mouse bite, and hole breakout [21, 48]. Annotation errors in the original dataset have been corrected, and representative samples are shown in Fig. 7.
Generation dataset. A subset of 5,252 defect images is reconstructed from DsPCBSD+ to train the diffusion model, as detailed in Tab. I. The training and validation sets are split at a ratio of 8:2.
Detection dataset. The remaining images are reorganized into a base detection set of 2,027 images (Tab. I, column “Base”), with all defects manually annotated using rectangular bounding boxes and split 8:2 into training and validation sets. To investigate the effect of different data expansion strategies, two extended variants are constructed:
  • Extend I (3,319 images): the base set expanded via conventional augmentations, including random flipping, rotation, and Gaussian blur.
  • Extend II (3,382 images): the base set expanded by incorporating defect images synthesized by the trained diffusion model. Annotations are produced semi-automatically: a seed set of real images is manually labeled, an auxiliary annotation model is trained on this seed set to label the synthetic images, and a final manual verification step corrects residual errors.
The two extended datasets are used in Sec. IV-C to provide a controlled comparison between traditional and generation-based augmentation strategies.
IV-B Experimental Setup
IV-B1 Implementation Details
The experimental environment in this study runs on an Ubuntu system, and the experimental setting employs an Intel Xeon E5-2686 v4 CPU and four 24G NVIDIA TITAN RTX GPUs as the hardware configuration. The generation model is based on Stable Diffusion v1.5 and uses a batch size of 1, 30,000 training steps, and a learning rate of 0.00002, while the detection model adopts a batch size of 4, is trained for 100 epochs, and uses a learning rate of 0.0001. For both the generation and detection tasks, the input image resolution is set to 512×512. All networks are optimized using the AdamW optimizer, where the momentum and weight decay are set to 0.9 and 0.0001, respectively.
IV-B2 Evaluation Metrics
The quality of synthesized PCB defect images is evaluated from three complementary perspectives. Fréchet Inception Distance (FID) [15] measures the distributional similarity between real and generated images in the feature space of a pretrained Inception network, where lower values indicate closer distributions:
$\mathrm{FID}=\|\mu_{r}-\mu_{g}\|^{2}+\mathrm{Tr}\!\left(\Sigma_{r}+\Sigma_{g}-2\left(\Sigma_{r}\Sigma_{g}\right)^{\frac{1}{2}}\right),$ (11)
where $\mu_{r},\Sigma_{r}$ and $\mu_{g},\Sigma_{g}$ are the mean and covariance of the real and generated feature distributions, respectively. Learned Perceptual Image Patch Similarity (LPIPS) [64] quantifies perceptual similarity by comparing deep feature activations across $L$ layers weighted by $w_{l}$, with lower values indicating greater perceptual fidelity:
$\mathrm{LPIPS}=\sum_{l}w_{l}\cdot\|\phi_{l}(x)-\phi_{l}(y)\|_{2}^{2}.$ (12)
Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) [55] assess pixel-level reconstruction quality and structural consistency, respectively, with higher values indicating better fidelity:
$\mathrm{PSNR}=10\cdot\log_{10}\!\left(\frac{\mathrm{MAX}^{2}}{\mathrm{MSE}}\right),$ (13)
$\mathrm{SSIM}=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (14)
where $\mu_{x},\mu_{y}$ are local means, $\sigma_{x}^{2},\sigma_{y}^{2}$ are variances, $\sigma_{xy}$ is the cross-covariance, and $C_{1},C_{2}$ are stabilizing constants.
Detection metrics. The downstream defect detection performance is evaluated using Precision, Recall, and mean Average Precision at two IoU thresholds (mAP@0.5 and mAP@0.5:0.95). Precision measures the fraction of correct detections among all detections, Recall measures the fraction of ground-truth defects successfully detected, and mAP summarizes the area under the precision–recall curve averaged over $N$ defect categories:
$\mathrm{Precision}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}},\quad\mathrm{Recall}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$ (15)
$\mathrm{mAP}=\frac{1}{N}\sum_{c=1}^{N}\int_{0}^{1}P(R)\,\mathrm{d}R.$ (16)
TABLE I: Dataset statistics for defect generation and detection experiments. Extend I: traditional augmentation; Extend II: generation-based augmentation.
Type Class Defect Gen. Defect Det.
Base Extend I Extend II
No. of Images Short 180 137 217 579
Spur 1,326 444 705 597
Spurious copper 826 360 597 613
Open 770 346 564 768
Mouse bite 1,084 422 705 592
Hole breakout 1,066 433 715 540
Total 5,252 2,027 3,319 3,382
No. of Defects Short 228 161 253 839
Spur 2,213 546 874 747
Spurious copper 969 385 641 667
Open 919 362 594 936
Mouse bite 1,309 440 732 636
Hole breakout 2,342 480 796 615
Total 7,980 2,374 3,890 4,440
IV-C Comparative Experiment
IV-C1 Detection Performance Comparison
To comprehensively evaluate UniPCB, we compare it against ten representative detectors spanning two architectural families. The Transformer-based group includes DETR [4], Deformable-DETR [70], DAB-DETR [30], DINO [62], RT-DETR [68], D-FINE [39], and DEIM [20]; the CNN-based group includes YOLOv8 [23], YOLOv10 [52], and YOLO11 [24]. RT-DETR serves as the direct baseline, as UniPCB is built upon it.
Quantitative results are reported in Tab. II, and visual comparisons are shown in Fig. 8. UniPCB achieves Precision / Recall / mAP@0.5 / mAP@0.5:0.95 of 0.977 / 0.953 / 0.980 / 0.618, outperforming all compared methods. Relative to the RT-DETR baseline, mAP@0.5 improves by 2.1% and mAP@0.5:0.95 by 1.7%, demonstrating that the IRSA and CLCF modules provide consistent gains beyond the base architecture. Among Transformer-based methods, the strongest competitor is DEIM (mAP@0.5: 0.958), which UniPCB surpasses by 2.2%. Among CNN-based methods, YOLOv8 achieves the best mAP@0.5:0.95 of 0.611, still 0.7% below UniPCB. The visual results in Fig. 8 further confirm that UniPCB produces fewer missed detections and tighter bounding boxes on small, low-contrast defects, where most competing methods show visible localization errors or omissions. The consistent improvement over all baselines trained on the same Extend II dataset suggests that the architectural gains from IRSA and CAMF are orthogonal to the data augmentation benefit, and that the two components of UniPCB are complementary.
TABLE II: Detection Performance of Different Models. The Top Three Results are Marked with red, blue, and green, respectively.
[…]
DETR [4] 20‘ECCV 0.896 0.875 0.923 0.582 66.6 60.5 41.5
Deformable-DETR [70] 21’ICLR 0.915 0.905 0.940 0.592 32.5 125.9 40.1
DAB-DETR [30] 22’ICLR 0.872 0.889 0.905 0.570 41.0 65.3 43.7
DINO [62] 23’ICLR 0.922 0.913 0.948 0.598 23.5 183.3 47.5
YOLOv8 [23] 23’Ultralytics 0.954 0.941 0.965 0.611 192.8 23.4 9.83
YOLOv10 [52] 24’NIPS 0.850 0.855 0.911 0.583 168.5 24.5 8.04
RT-DETR [68] 24’CVPR 0.949 0.921 0.959 0.601 97.0 57.2 19.9
YOLO11 [24] 24’Ultralytics 0.956 0.921 0.960 0.607 192.3 21.3 9.42
D-FINE [39] 25’ICLR 0.938 0.927 0.952 0.606 150.6 56.3 19.1
DEIM [20] 25’CVPR 0.942 0.938 0.958 0.610 118.2 92.5 31.3
UniPCB (Detection) - 0.977 0.953 0.980 0.618 90.9 71.8 19.1
Figure 8: Visualization of detection results across different models. Zoom-in for best view.
IV-C2 Generation Performance Comparison
We compare UniPCB against three representative conditional generation baselines: ControlNet [63], which injects a single condition at the initial resolution; Uni-ControlNet [67], which combines edge and depth conditions through a unified control mechanism but still injects them at a single scale; and AnyControl [45], which supports flexible multi-condition combinations but without explicit cross-modal fusion. To further isolate the effect of condition type, we also evaluate a variant UniPCB (+seg) that replaces the text condition with a defect segmentation mask as the third control signal. Results are reported in Tab. IV and visualized in Fig. 9.
The results reveal a clear progression. ControlNet, relying solely on edge maps, achieves PSNR/SSIM of 10.23/0.502 but yields a relatively high FID of 131.11, as single-modality conditions cannot fully constrain the structural diversity of PCB layouts. Uni-ControlNet improves PSNR to 10.60 by adding depth as a complementary condition, yet its FID rises to 145.53 and SSIM drops to 0.457, suggesting that naïve concatenation of conditions without dedicated fusion introduces inter-modal interference. AnyControl shows similar limitations despite its flexible condition combinations, achieving the weakest overall performance (FID: 142.01, SSIM: 0.404). UniPCB (+seg) improves PSNR and SSIM over the baselines by adding a segmentation mask, but its FID (142.93) and LPIPS (0.535) remain worse than UniPCB, indicating that a pixel-level mask introduces spatial redundancy with the edge condition rather than complementary information.
UniPCB achieves the best performance across all metrics (FID: 129.61, LPIPS: 0.457, PSNR: 13.45, SSIM: 0.619). The key distinction from prior methods is not the number of conditions but their complementarity and fusion quality: edge, depth, and text provide structural, spatial, and semantic cues that cover orthogonal axes of control, while ScaleEncoder and CondMod ensure that these conditions are effectively embedded across multiple scales rather than injected at a single resolution. The qualitative results in Fig. 9 further confirm that UniPCB preserves the global circuit topology while rendering localized defect patterns with higher fidelity than all compared methods.
TABLE III: Evaluation of augmentation strategies. Superscripts $\mathrm{I}$ and $\mathrm{II}$ denote traditional augmentation and generative augmentation, respectively. For results obtained with generative augmentation, the percentages in parentheses indicate relative changes compared with the corresponding traditional augmentation setting. red indicates favorable changes, while green indicates unfavorable changes.
[…]
DETRI [4] 20’ECCV 0.892 0.848 0.914 0.556
DETRII [4] 0.896 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.4\%$} 0.875 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.7\%$} 0.923 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.9\%$} 0.582 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.6\%$}
Deformable-DETRI [70] 21’ICLR 0.917 0.882 0.932 0.561
Deformable-DETRII [70] 0.915 $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\,\downarrow 0.2\%$} 0.905 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.3\%$} 0.940 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.8\%$} 0.592 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 3.1\%$}
DAB-DETRI [30] 22’ICLR 0.869 0.867 0.899 0.545
DAB-DETRII [30] 0.872 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.3\%$} 0.889 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.2\%$} 0.905 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.6\%$} 0.570 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.5\%$}
DINOI [62] 23’ICLR 0.928 0.888 0.940 0.574
DINOII [62] 0.922 $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\,\downarrow 0.6\%$} 0.913 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.5\%$} 0.948 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.8\%$} 0.598 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.4\%$}
YOLOv8I [23] 23’Ultralytics 0.945 0.933 0.956 0.601
YOLOv8II [23] 0.954 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.9\%$} 0.941 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.8\%$} 0.965 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.9\%$} 0.611 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 1.0\%$}
YOLOv10I [52] 24’NIPS 0.853 0.840 0.901 0.564
YOLOv10II [52] 0.850 $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\,\downarrow 0.3\%$} 0.855 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 1.5\%$} 0.911 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 1.0\%$} 0.583 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 1.9\%$}
RT-DETRI [68] 24’CVPR 0.951 0.895 0.950 0.573
RT-DETRII [68] 0.949 $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\,\downarrow 0.3\%$} 0.921 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.6\%$} 0.959 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.9\%$} 0.601 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.8\%$}
YOLO11I [24] 24’Ultralytics 0.927 0.899 0.952 0.574
YOLO11II [24] 0.956 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.9\%$} 0.921 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.2\%$} 0.960 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.8\%$} 0.607 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 3.3\%$}
D-FINEI [39] 25’ICLR 0.935 0.910 0.944 0.579
D-FINEII [39] 0.938 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.3\%$} 0.927 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 1.7\%$} 0.952 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.8\%$} 0.606 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.7\%$}
DEIMI [20] 25’CVPR 0.940 0.914 0.948 0.584
DEIMII [20] 0.942 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.3\%$} 0.938 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.4\%$} 0.958 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 1.0\%$} 0.610 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.6\%$}
UniPCBI - 0.970 0.927 0.971 0.593
UniPCBII 0.977 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.7\%$} 0.953 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.6\%$} 0.980 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 0.9\%$} 0.618 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,\uparrow 2.1\%$}
TABLE IV: Generation Performance of Different Models.
[…]
ControlNet [63] 131.11 0.566 10.23 0.502
Uni-Controlnet [67] 145.53 0.562 10.60 0.457
AnyControl [45] 142.01 0.635 9.02 0.404
UniPCB (+seg) 142.93 0.535 12.47 0.525
UniPCB (ours) 129.61 0.457 13.45 0.619
Figure 9: Visualization of generation results across different models. For better clarity, results are best viewed in enlarged form.
IV-C3 Data Augmentation Comparison
To evaluate the downstream benefit of generation-based augmentation, all detection models in Tab. II are retrained on Extend I and Extend II respectively, and their performance is systematically compared in Tab. III.
Models trained on Extend II consistently outperform those trained on Extend I across almost all metrics. The gains are most pronounced in Recall and mAP@0.5:0.95, where most models improve by 1.5%–2.0%, reflecting that generation-augmented data helps reduce missed detections and improves localization quality at stricter IoU thresholds. mAP@0.5 also shows consistent improvement close to 1%. Precision remains nearly unchanged or shows slight declines for some models; the minor degradation is likely attributable to additional false positives caused by enhanced sensitivity to defect features introduced by the richer training distribution.
These results confirm that the generation-based augmentation provides richer and more informative training samples than conventional augmentation, enabling detection models to better capture discriminative features across all six defect categories. Notably, UniPCB itself benefits from Extend II with improvements of ++0.7% / ++2.6% / ++0.9% / ++2.1% in Precision / Recall / mAP@0.5 / mAP@0.5:0.95, demonstrating that the generation and detection components of UniPCB are mutually reinforcing.
Figure 10: Examples of synthesized defect images with bounding-box annotations, covering all six categories in Fig. 7. Annotations are generated by an auxiliary model trained on seed annotations and verified manually.
TABLE V: Ablation Study on The Effectiveness of IRSA and CLCF Modules.
[…]
Baseline 0.949 0.921 0.959 0.601
a 0.951 0.933 0.968 0.614
b 0.969 0.949 0.969 0.614
c 0.977 0.953 0.980 0.618
TABLE VI: Ablation results of different attention mechanisms in CLCF Block.
[…]
Baseline 0.949 0.921 0.959 0.601
+SE [18] 0.945 0.932 0.960 0.604
+CA [17] 0.936 0.908 0.945 0.582
+SimAM [58] 0.958 0.918 0.965 0.607
+EMA [37] 0.963 0.913 0.964 0.609
+DPCA (ours) 0.969 0.949 0.969 0.614
IV-D Ablation Study
IV-D1 Module contribution
Tab. V reports the contribution of each proposed module. Adding IRSA alone (setting (a)) yields gains of ++0.9% / ++1.3% in mAP@0.5 / mAP@0.5:0.95 over the baseline, with a notable Recall improvement of ++1.2%. This suggests that coupling self-attention with shift-wise convolution reduces missed detections by capturing fine-grained local cues that vanilla self-attention overlooks, particularly for small, low-contrast defects embedded in complex circuit patterns. Adding CLCF alone (setting (b)) produces larger Precision gains (++2.0%) alongside comparable mAP improvements, indicating that the pixel-level gating mechanism effectively suppresses false detections caused by background interference during feature fusion. When both modules are combined (setting (c)), the model achieves the best performance across all metrics, with mAP@0.5 and mAP@0.5:0.95 reaching 0.980 and 0.618, improving the baseline by ++2.1% and ++1.7% respectively. The complementary nature of the two modules is evident: IRSA strengthens single-scale feature discrimination while CLCF enhances cross-level feature aggregation, and their combination yields gains exceeding either module alone.
IV-D2 Attention mechanism comparison
To validate the design of the DPCA sub-block within CLCF, we replace it with four commonly used attention mechanisms and report results in Tab. VI.
SE improves Recall by ++1.1% through inter-channel reweighting but lacks spatial modeling, limiting overall gains. CA introduces positional encoding to enhance spatial awareness yet shows degraded performance across all metrics (mAP@0.5:0.95 drops by 1.9% relative to baseline), indicating poor compatibility with the cross-level fusion objective. SimAM achieves moderate and consistent gains (++0.6% mAP@0.5:0.95) without additional parameters. EMA further improves through cross-scale interactions, reaching Precision / mAP@0.5:0.95 of 0.963 / 0.609.
DPCA outperforms all alternatives across every metric, with Precision / Recall / mAP@0.5 / mAP@0.5:0.95 of 0.969 / 0.949 / 0.969 / 0.614 (++2.0% / ++2.8% / ++1.0% / ++1.3% over baseline). The gains are attributable to its parallel local–global design: the local branch captures fine-grained spatial cues via grouped convolutions, while the global branch models long-range dependencies via scaled self-attention, and their combination generates spatially adaptive weight maps that are better suited to the diverse scales and locations of PCB defects than any single-path mechanism.
V Conclusion and Limitation
In this paper, we present UniPCB, a generation-assisted PCB defect inspection pipeline that integrates controlled defect synthesis with task-specific defect detection. To alleviate data scarcity and class imbalance, we design a Multi-Modal Condition Generator to derive complementary edge, depth, and text conditions from PCB images. Based on these conditions, we introduce a diffusion-based synthesis network with a ScaleEncoder and Conditional Modulation Blocks, which inject multi-resolution control into the denoising process and generate structurally aligned defect samples. To improve defect representation under complex circuit backgrounds, we further develop an Inverted Residual Shift Attention Block for joint local–global feature modeling and a Cross-level Complementary Fusion Block for selective cross-level feature refinement. Experiments on DsPCBSD+ show that the proposed generation-assisted pipeline improves both synthetic sample quality and downstream detection performance over the compared baselines.
Despite these results, UniPCB does not perform end-to-end joint optimization between the generator and detector. Instead, it provides a coordinated pipeline in which the generator supplies structurally controlled defect samples and the detector is designed to exploit the enriched training data. In addition, the current evaluation is mainly conducted on DsPCBSD+-derived datasets, and further validation on PCB data collected from different AOI devices, production lines, and board layouts is needed to assess cross-domain robustness. Exploring detector-feedback-driven generation and broader industrial deployment scenarios remains future work.

### 26 May 2026 v3

@@
 Abstract
-Printed Circuit Board (PCB) defect inspection faces two compounding challenges: scarce and imbalanced defect samples that limit model training, and insufficient feature representation under complex circuit backgrounds.
-Existing generation methods rely on single-modality conditions with coarse structural control, while detection methods improve architectures without addressing the data bottleneck.
-To resolve both challenges jointly, we propose a generation-assisted PCB defect inspection framework that integrates controlled defect synthesis with task-specific defect detection.
+Accurate and efficient printed circuit board (PCB) defect inspection is critical for ensuring product reliability in automated electronics manufacturing.
+However, PCB defect inspection faces two compounding challenges: scarce and imbalanced defect samples that limit model training, and insufficient feature representation under complex circuit backgrounds.
+Existing generation methods often rely on single-modality conditions with coarse structural control, while detection methods mainly improve model architectures without explicitly addressing the data bottleneck.
+To address these challenges, we propose UniPCB, a generation-assisted PCB defect inspection framework that integrates controlled defect synthesis with task-specific defect detection.
 On the generation side, a Multi-modal Condition Generator extracts complementary edge, depth, and text conditions in parallel.
-A ScaleEncoder then embeds these conditions into the diffusion U-Net at four resolutions, and a Condition Modulation applies FiLM-style spatially-adaptive modulation at each scale, enabling structurally aligned and defect-aware sample synthesis.
-On the detection side, an Inverted Residual Shift Attention couples self-attention with shift-wise convolution to jointly capture global context and local texture, and a Cross-level Complementary Fusion Block generates pixel-level gates for selective cross-level feature fusion.
-The synthesized samples directly enrich the detection training set, so that improvements in generation compound with improvements in detection.
-Extensive experiments on DsPCBSD+ demonstrate that UniPCB achieves mAP@0.5 of 98.0% and mAP@0.5:0.95 of 61.8% on defect detection, surpassing all compared methods, while the generation branch attains an FID of 129.61 and SSIM of 0.619, outperforming existing conditional generation approaches.
+A ScaleEncoder embeds these conditions into the diffusion U-Net at four resolutions, and a Condition Modulation block applies FiLM-style spatially adaptive modulation at each scale, enabling structurally aligned and defect-aware sample synthesis.
+On the detection side, an Inverted Residual Shift Attention block couples self-attention with shift-wise convolution to jointly capture global context and local texture, while a Cross-level Complementary Fusion block generates pixel-level gates for selective cross-level feature fusion.
+The synthesized samples enrich the detection training set, allowing the detector to benefit from more diverse and informative defect observations.
+Extensive experiments on DsPCBSD+ demonstrate that UniPCB achieves an mAP@0.5 of 98.0% and an mAP@0.5:0.95 of 61.8% on defect detection, outperforming the compared methods.
+The generation branch also achieves an FID of 129.61 and an SSIM of 0.619, showing improved conditional defect synthesis quality over existing baselines.
+Code is available at: https://github.com/House-yuyu/UniPCB.
 Index Terms:
-PCB defect detection, Diffusion model, Multi-modal data generation.
+Printed circuit board defect detection, conditional diffusion model, generation-assisted inspection.
 I Introduction
-Printed circuit boards (PCBs) are essential to modern electronics, but complex manufacturing processes can introduce defects such as shorts, opens, and spurious copper that threaten product reliability.
-Automated optical inspection (AOI) systems typically rely on reference-image comparison and handcrafted feature matching, making them sensitive to lighting variation, board layout changes, and novel defect morphologies, which can lead to high false-negative rates in production environments [69, 14].
-Deep learning-based detectors offer substantially greater robustness by learning discriminative representations directly from data.
-However, their effectiveness depends on the availability of large volumes of labeled defect samples, which is fundamentally difficult to satisfy in PCB manufacturing.
-Defective boards are rare by design, and defect categories often follow a severely long-tailed distribution.
-As a result, models trained on limited available data tend to underfit rare defect categories and generalize poorly to unseen board layouts.
-To alleviate this data bottleneck, traditional augmentation techniques such as rotation and flipping have been widely adopted, but they fail to capture the complex and variable defect patterns encountered in real production environments.
-Deep generative models offer a more principled alternative: VAEs [25, 51] and GANs [11] have been applied to defect synthesis, but VAEs tend to produce blurry images due to pixel-wise reconstruction losses, and GANs suffer from training instability and mode collapse.
-Diffusion models [16, 43, 44, 66, 42, 65] overcome these limitations with superior sample quality and diversity.
-Nonetheless, their training on natural image datasets limits their effectiveness on industrial images such as PCBs [5, 13], and existing PCB-oriented methods typically rely on a single control condition injected at the initial resolution, providing insufficient structural guidance for the dense, fine-grained circuit patterns unique to PCB images.
+Printed circuit boards (PCBs) are essential components in modern electronic products, and their manufacturing quality directly affects the reliability of downstream devices.
+Complex fabrication processes may introduce defects such as shorts, opens, mouse bites, spurs, and spurious copper, which can degrade electrical connectivity and cause product failure.
+Automated optical inspection (AOI) systems are widely used for PCB quality control, but traditional reference-image comparison and handcrafted feature matching are sensitive to lighting variations, board layout changes, and novel defect morphologies, often leading to missed detections or false alarms in practical inspection scenarios [1, 2].
+Deep learning-based detectors improve robustness by learning discriminative representations directly from data.
+However, their effectiveness depends on large and well-annotated defect datasets, which are difficult to obtain in PCB manufacturing.
+First, mature production lines usually have high yield rates, making defective samples naturally rare.
+Second, PCB layouts vary across products, limiting the transferability of defect samples across board types.
+Third, defect annotation requires domain expertise and is time-consuming, especially when dense and small defects need to be precisely localized.
+As a result, PCB defect datasets are often scarce and class-imbalanced, causing detectors to underfit rare defects and generalize poorly to unseen board layouts.
+To alleviate this data bottleneck, traditional augmentation techniques such as rotation, flipping, and blurring have been widely adopted.
+Nevertheless, these operations mainly transform existing samples and cannot sufficiently model the diverse defect morphologies and local structural variations observed in real PCB images.
+Deep generative models provide a more flexible alternative.
+VAEs [3, 4] and GANs [5] have been explored for defect synthesis, but VAEs often produce blurry details and GANs may suffer from training instability and mode collapse.
+Diffusion models [6, 7, 8, 9, 10, 11] offer stronger training stability and better texture modeling ability, making them suitable for fine-grained industrial defect synthesis.
+However, their training on natural image datasets limits their effectiveness on industrial images such as PCBs [12, 13], and existing PCB-oriented methods typically rely on a single control condition injected at the initial resolution, providing insufficient structural guidance for the dense, fine-grained circuit patterns unique to PCB images.
 Figure 1: Comparison between UniPCB and previous approaches for PCB defect detection.
@@
 (b) UniPCB jointly addresses both within a unified framework: the generation branch synthesizes high-quality defect samples while the detection branch exploits the enriched data through task-specific feature modeling modules.
-Beyond high-quality sample generation, the expressive capability of the detection framework itself is equally critical.
-DETR [4] leverages end-to-end global modeling to significantly enhance defect perception in complex scenes, but its slow convergence and high computational cost hinder real-time industrial deployment.
-RT-DETR [68] improves inference speed while maintaining competitive accuracy through optimized attention mechanisms and detection head design, making it a more practical alternative.
-However, under complex backgrounds or small defects, its feature modeling and multi-scale feature fusion remain insufficient.
-The challenge of simultaneously optimizing detection accuracy and computational efficiency has been extensively studied in the general detection literature [47], yet achieving this balance for fine-grained PCB defects under dense circuit patterns remains an open problem [50].
+In addition to data scarcity, the feature representation capability of the detector remains critical.
+Transformer-based detectors such as DETR [14] and RT-DETR [15] provide strong global modeling ability, while lightweight and efficient detection designs have been widely studied to balance accuracy and computational cost [16, 17].
+Recent studies in surface defect and tiny object detection further indicate that fine-grained feature separation, relation reasoning, class knowledge, and regional attention are important for detecting subtle defects under complex backgrounds [18, 19, 20].
+Nevertheless, architecture-level improvements alone may still be constrained by insufficient and imbalanced training data [21].
 These two issues often appear together in PCB inspection and jointly limit practical detection performance.
@@
 Traditional handcrafted or template-matching methods are sensitive to manufacturing variations and therefore generalize poorly across board layouts and imaging conditions.
-Deep learning based detectors have improved robustness by learning discriminative representations automatically, and recent studies [60, 54, 59, 26, 32, 53, 34] mainly improve PCB defect detection from two perspectives: adaptive feature modeling and multi-scale feature fusion.
-Adaptive feature modeling [31, 28, 12] aims to enhance defect-relevant responses while suppressing complex circuit background interference.
+Deep learning based detectors have improved robustness by learning discriminative representations automatically, and recent PCB defect detection studies [32, 33, 34, 35, 36] mainly improve performance from two perspectives: adaptive feature modeling and multi-scale feature fusion.
+Beyond PCB-specific scenarios, recent TCSVT studies on industrial defect and anomaly detection have also explored knowledge distillation for rail surface defects [37], differential distillation for unsupervised anomaly localization [38], multi-perspective multimodal fusion for 3D industrial defects [39], and memory-augmented wavelet representations for anomaly detection [40].
+These works further indicate that efficient representation learning and robust localization are central issues in industrial visual inspection.
+Adaptive feature modeling [41, 42] aims to enhance defect-relevant responses while suppressing complex circuit background interference.
 Liu et al.
@@
 [46] introduced a Slim-Scale Adaptive Fusion structure that aggregates hierarchical features using grouped spatial convolutional pyramids and channel shuffle, improving fusion efficiency and detection accuracy.
-Although these methods improve feature representation and localization, they mainly operate at the detector architecture level and typically assume that the training distribution is sufficiently representative.
+Although these PCB-specific and general industrial inspection methods improve feature representation, localization, and deployment efficiency, they mainly operate at the detector or anomaly modeling level and typically assume that the training distribution is sufficiently representative.
 In practical PCB inspection, however, defect samples remain scarce and highly imbalanced, especially for rare defect categories.
@@
 Figure 9: Visualization of generation results across different models. For better clarity, results are best viewed in enlarged form.
-IV-C3 Data Augmentation Comparison
-To evaluate the downstream benefit of generation-based augmentation, all detection models in Tab.
-II are retrained on Extend I and Extend II respectively, and their performance is systematically compared in Tab.
-III.
-Models trained on Extend II consistently outperform those trained on Extend I across almost all metrics.
-The gains are most pronounced in Recall and mAP@0.5:0.95, where most models improve by 1.5%–2.0%, reflecting that generation-augmented data helps reduce missed detections and improves localization quality at stricter IoU thresholds.
-mAP@0.5 also shows consistent improvement close to 1%.
-Precision remains nearly unchanged or shows slight declines for some models; the minor degradation is likely attributable to additional false positives caused by enhanced sensitivity to defect features introduced by the richer training distribution.
-These results confirm that the generation-based augmentation provides richer and more informative training samples than conventional augmentation, enabling detection models to better capture discriminative features across all six defect categories.
-Notably, UniPCB itself benefits from Extend II with improvements of ++0.7% / ++2.6% / ++0.9% / ++2.1% in Precision / Recall / mAP@0.5 / mAP@0.5:0.95, demonstrating that the generation and detection components of UniPCB are mutually reinforcing.
-Figure 10: Examples of synthesized defect images with bounding-box annotations, covering all six categories in Fig.
-7.
-Annotations are generated by an auxiliary model trained on seed annotations and verified manually.
 TABLE V: Ablation Study on The Effectiveness of IRSA and CLCF Modules.
@@
 +DPCA (ours) 0.969 0.949 0.969 0.614
+IV-C3 Data Augmentation Comparison
+To evaluate the downstream benefit of generation-based augmentation, all detection models in Tab.
+II are retrained on Extend I and Extend II respectively, and their performance is systematically compared in Tab.
+III.
+Models trained on Extend II consistently outperform those trained on Extend I across almost all metrics.
+The gains are most pronounced in Recall and mAP@0.5:0.95, where most models improve by 1.5%–2.0%, reflecting that generation-augmented data helps reduce missed detections and improves localization quality at stricter IoU thresholds.
+mAP@0.5 also shows consistent improvement close to 1%.
+Precision remains nearly unchanged or shows slight declines for some models; the minor degradation is likely attributable to additional false positives caused by enhanced sensitivity to defect features introduced by the richer training distribution.
+These results confirm that the generation-based augmentation provides richer and more informative training samples than conventional augmentation, enabling detection models to better capture discriminative features across all six defect categories.
+Notably, UniPCB itself benefits from Extend II with improvements of ++0.7% / ++2.6% / ++0.9% / ++2.1% in Precision / Recall / mAP@0.5 / mAP@0.5:0.95, demonstrating that the generation and detection components of UniPCB are mutually reinforcing.
+Figure 10: Examples of synthesized defect images with bounding-box annotations, covering all six categories in Fig.
+7.
+Annotations are generated by an auxiliary model trained on seed annotations and verified manually.
 IV-D Ablation Study
@@
 V Conclusion and Limitation
-In this paper, we present UniPCB, a generation-assisted PCB defect inspection pipeline that integrates controlled defect synthesis with task-specific defect detection.
+In this paper, we present UniPCB, a generation-assisted PCB defect inspection pipeline that integrates controlled defect synthesis with task-specific defect detection pipeline.
 To alleviate data scarcity and class imbalance, we design a Multi-Modal Condition Generator to derive complementary edge, depth, and text conditions from PCB images.
@@
 Instead, it provides a coordinated pipeline in which the generator supplies structurally controlled defect samples and the detector is designed to exploit the enriched training data.
-In addition, the current evaluation is mainly conducted on DsPCBSD+-derived datasets, and further validation on PCB data collected from different AOI devices, production lines, and board layouts is needed to assess cross-domain robustness.
+In addition, the current evaluation is mainly conducted on DsPCBSD+-derived datasets, and further validation on PCB data collected from different production lines, AOI devices, and board layouts is needed to assess cross-domain robustness.
 Exploring detector-feedback-driven generation and broader industrial deployment scenarios remains future work.